# Multi-Model Synthetic Training for Mission-Critical Small Language Models

### v1

@@ Abstract
Index Terms:
natural language processing, large language models, small language models, domain-specific fine tuning, supervised learning, supervised fine-tuning (SFT), cost efficient quantization, domain-specific SLM experts
@@ III-D2 Manual Evaluation
To validate our automatic pipeline and assess qualitative aspects, we manually evaluated 100 randomly sampled responses. Manual evaluation followed identical numerical accuracy criteria but additionally assessed:
  • Correctness of reasoning steps
  • Presence of calculation errors despite correct final answers
  • Quality of explanations and maritime domain understanding
@@ V Results
Our evaluation reveals that traditional NLP metrics fail to show language models’ performance in very specialized domains. As shown in Table III, despite having near-zero BLEU scores, manual evaluation underlines the model has strong accuracy and consistency.

### v2

@@ Abstract
**footnotetext: Work done during an internship at Hitachi Vantara Federal.
@@ III-B Synthetic Q&A Generation Pipeline
Example queries include: ”Which vessels near Port of Los Angeles changed course by more than 45 degrees in the past hour?” (Movement Analysis), and ”Identify any cargo ships exceeding 25 knots within 5 nautical miles of San Francisco Bay” (Anomaly Detection).
@@ III-D2 Manual Evaluation
To validate our automatic pipeline and assess qualitative aspects, we manually evaluated 100 randomly sampled responses. Manual evaluation followed identical numerical accuracy criteria but additionally assessed the correctness of reasoning steps, presence of calculation errors despite correct final answers, and the quality of explanations and maritime domain understanding.
@@ V Results
Our evaluation reveals that traditional NLP metrics fail to show language models’ performance in very specialized domains. As shown in Table III, despite having near-zero BLEU scores, manual evaluation underlines the model [31] has strong accuracy and consistency.